# The Resolution of the Gauge Problem of Cosmology provides Insight into the Formation of the First Structures in the Universe

### v2

@@ III Gauge Problem of Cosmology
Einstein’s equations and conservation laws are invariant under general coordinate transformations $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$. This signifies that there are no preferred coordinate systems. Consequently, the linearized Einstein equations and conservation laws are invariant under a general linear infinitesimal space-time transformation, which is usually referred to as a gauge transformation. This transformation is given by
@@
where the gauge functions $\xi^{\mu}(t,\bm{x})$ are four arbitrary infinitesimal functions of time, $x^{0}=ct$ with $c$ the speed of light, and space, $\bm{x}=(x^{1},x^{2},x^{3})$, coordinates.
@@
where $u^{\mu}$ is the fluid four-velocity normalized to unity, ${u^{\mu}u_{\mu}=1}$, and $N^{\mu}:=nu^{\mu}$ represents the cosmological particle current four-vector, which satisfies the particle number conservation law ${N^{\mu}{}_{;\mu}=0}$. A semicolon denotes covariant differentiation.
@@
Let $S_{\scriptscriptstyle(0)}(t)$ denote the quantities $\varepsilon_{\scriptscriptstyle(0)}$, $n_{\scriptscriptstyle(0)}$, and $\theta_{\scriptscriptstyle(0)}$. These quantities satisfy the background, that is to say, the unperturbed Einstein equations and conservation laws. Let $S_{\scriptscriptstyle(1)}(t,\bm{x})$ be their perturbed counterparts $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, which satisfy the linearized equations. As a consequence of the linearity of these equations, new solutions that are physically equivalent can be generated. These solutions are given by (see Weinberg [31], Sec. 10.9 for a detailed explanation)
$S^{\prime}_{\scriptscriptstyle(1)}=S_{\scriptscriptstyle(1)}+\mathcal{L}_{\xi}S_{\scriptscriptstyle(0)}=S_{\scriptscriptstyle(1)}+\xi^{0}\dot{S}_{\scriptscriptstyle(0)}.$ (3)
@@
In this expression, the operator $\mathcal{L}_{\xi}$ is the Lie derivative with respect to the infinitesimal four-vector $\xi^{\mu}$. An overdot denotes differentiation with respect to $x^{0}=ct$. The solutions $S_{\scriptscriptstyle(1)}$ and $S^{\prime}_{\scriptscriptstyle(1)}$ contain the so-called gauge modes which are given by $\hat{S}_{\scriptscriptstyle(1)}=\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$. This indicates that the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ are dependent on the choice of a system of reference, rendering them gauge-dependent and, consequently, devoid of physical significance. This is the gauge problem of cosmology, whereby the coordinate artifacts and the underlying physics are inextricably linked in the solution of the linearized equations. Even when physical initial values are imposed in order to solve the linearized equations, the resulting solution lacks physical meaning. It is therefore imperative to resolve the gauge problem in order to facilitate an investigation into the evolution of density perturbations in the universe.
IV Solving the Gauge Problem of Cosmology
@@
In contrast to the gauge-dependent quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, their physical counterparts are independent of the choice of a coordinate system. This is to say that they are gauge-invariant. As the physics of the density perturbations is concealed within the general solution of the linearized Einstein equations and conservation laws, it follows that the physical counterparts of $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, must be linear combinations of gauge-dependent solutions to these equations. In order to identify these quantities, it is necessary to consider the fact that a quantity is gauge-invariant if it is independent of the coordinate transformation parameter $\xi^{\mu}$, as outlined by Stewart [3]. This implies that the quantity must be invariant under a linear coordinate transformation (1). As is evident from (3), the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ are independent of the spatial components, $\xi^{i}$, of the transformation, but not of the temporal component, $\xi^{0}$. Given that the transformation of these quantities under the general infinitesimal transformation (1) is identical to that described by (3), it is possible to construct gauge-invariant quantities by combining these gauge-dependent quantities in a way that eliminates the gauge modes $\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$. This results in three distinct sets of linear combinations. In each of these sets, precisely one gauge-invariant quantity is identically zero. Given that the focus is on energy density and particle number density perturbations, the only possible set is
@@
Local density perturbations exert no influence on the expansion of the universe, as evidenced by (4b). The transformation rule (3), illustrates that the quantities (4) are indeed gauge-invariant.
In Sec. IX, it is demonstrated that in the nonrelativistic limit, where the pressure is considered to be zero, the linearized conservation laws lack physical significance and are decoupled from the linearized energy density constraint equation and momentum constraint equations. The combination of these constraint equations results in the time-independent Poisson equation of Newtonian gravitation with a source term $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})$. Furthermore, the special relativistic relation $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}mc^{2}$, where $m$ is the rest mass, is recovered. In light of the aforementioned considerations, it can be concluded that the quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent the actual physical perturbations of the energy density and the particle number density, respectively, in the context of general relativity.
The sole method for constructing gauge-invariant quantities for the energy density and particle number density perturbations is provided by the expressions (4). As these quantities also possess the correct nonrelativistic limit, it can be concluded that the gauge problem of cosmology has been resolved. This outcome is an inevitable result of the mathematical consistency of general relativity. That is to say, the general theory of relativity allows for only one gauge-invariant quantity that accurately represents a physical density perturbation.
V Selection of a Reference System
@@
The quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ are defined without the use of a coordinate system. However, to derive evolution equations for these quantities, it is first necessary to establish a system of reference. Since general relativity is covariant, and the quantities in question are gauge-invariant, it follows that the evolution equations for these quantities can be derived in any chosen coordinate system. The selection of an appropriate coordinate system for this problem is dependent on two criteria, which are outlined below.
@@
Firstly, the interpretation of $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ requires taking the nonrelativistic limit. In the context of general relativity, coordinate transformations are in general space-time transformations, specifically $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$. In the linear case, space-time transformations are given by (1). In Newtonian gravitation, the space and time transformations are regarded as distinct and independent entities. Consequently, when taking the nonrelativistic limit, the relativistic space-time transformations must be automatically separated into independent space and time transformations. Secondly, it would be advantageous to have a coordinate system that would facilitate the derivation of the evolution equations. It will now be demonstrated that both of these requirements can be met by employing the same coordinate system.
@@
In Newtonian gravitation, where space and time are treated as distinct and independent entities, all coordinate systems are inherently synchronous. In view of the nonrelativistic limit, it can be seen that a synchronous reference system represents the optimal choice. In these coordinates, the metric tensor $g_{\mu\nu}(t,\bm{x})$ of the flrw universes is given by
@@
The scale factor of the universe is designated as $a(t)$. Given that $g_{00}=1$, coordinate time is equal to proper time. The synchronicity condition is $g_{0i}=0$, as explained in Ref. [31], Sec. 11.8 and Ref. [35], § 84. The metric tensor of the three-dimensional maximally symmetric subspaces of constant time is denoted by $\tilde{g}_{ij}$. The Killing equations, $\xi_{\mu;\nu}+\xi_{\nu;\mu}=0$, and (5) demonstrate that the gauge functions $\xi^{\mu}(t,\bm{x})$ in the transformation (1) become
@@
if only transformations between synchronous coordinates are allowed. In expressions (6), $\psi(\bm{x})$ and $\chi^{i}(\bm{x})$ are four arbitrary infinitesimal functions of the spatial coordinates. In Sec. IX, it will be proven that in the nonrelativistic limit, the function $\psi(\bm{x})$ is an arbitrary infinitesimal constant. This means that the transformations of space and time are independent of each other in this limit.
The second requirement will now be considered. Synchronous coordinates possess the property that the space-space components, $R_{ij}$, of the four-dimensional Ricci curvature tensor, $R_{\mu\nu}$, are partitioned into two distinct parts. One part exclusively comprises the time derivatives of the spatial metric tensor, $g_{ij}$, while the second part is the Ricci curvature tensor of the three-dimensional subspaces, denoted by ${{}^{3}\!}R_{ij}$. This is demonstrated in Ref. [35], § 97. As the perturbations of the spatial metric tensor and the spatial Ricci tensor are also tensors (see Sec. VI.2), the covariant decomposition theorems in Refs. [34, 36, 3] may be employed to decompose these perturbed tensors into their scalar, vector, and tensor parts. This decomposition permits the system of linearized Einstein equations and conservation laws to be decomposed into three independent systems, each of which describes the evolution of scalar, vector, and tensor perturbations. As will be demonstrated in Sec. VII, only scalar perturbations are associated with density perturbations. Accordingly, it is sufficient to consider the system for scalar perturbations (37). The aforementioned system is more straightforward to utilize than the original set of equations (12), thus facilitating the derivation of the evolution equations for density fluctuations in synchronous coordinates in comparison to any other coordinate system. In view of the fact that synchronous coordinates are likewise compatible with Newton’s theory of gravitation, they will be employed in the derivation of the primary result (60).
@@ VI Einstein Equations and Conservation Laws
In the preceding section, it was concluded that synchronous coordinates are the optimal choice for addressing the problem at hand. In this section, the background Einstein equations and conservation laws, along with their linearized counterparts, will be expressed in this particular coordinate system.
The background energy density constraint equation and its linearized counterpart are expressed in a manner consistent with the contracted Bianchi identities. These identities state that the Riemann curvature tensor satisfies the equation $R^{\mu}{}_{\nu;\mu}=\tfrac{1}{2}R_{,\nu}$. This can be expressed in an equivalent manner as the four-divergence of the Einstein tensor, i.e., ${G^{\mu\nu}{}_{;\nu}=0}$. This result demonstrates that the constraint equations contain at most first-order time derivatives of the metric, which is precisely what is required for the development of the novel perturbation theory. For further details see Ref. [31], Secs. 6.8 and 7.5 and Ref. [35], § 92 and § 95.
It is understood from thermodynamic principles that the energy density, $\varepsilon$, and the pressure, $p$, are functions of both the particle number density, $n$, and the temperature, $T$:
$\varepsilon=\varepsilon(n,T),\quad p=p(n,T).$ (7)
$p=p(\varepsilon,n),$ (8)
which will be utilized in the subsequent analysis.
@@ VI.1 Background Equations
It can be demonstrated that the $G_{0i}$ constraint equations and the $G_{ij}$ dynamical equations with $i\neq j$ are satisfied identically. The $G_{ii}$ dynamical equations are equivalent to the time derivative of the $G_{00}$ constraint equation (Friedmann equation) given by equation (10a). Therefore, the $G_{ij}$ dynamical equations can be omitted. The cosmological constant, $\Lambda$, the gravitational constant, $G_{\text{N}}$, and the speed of light, $c$, are fundamental constants.
The overdot is used to denote differentiation with respect to $x^{0}=ct$. The Hubble function $H$ is defined by $H:=\dot{a}/a$. For flrw universes, the Hubble function is given by $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, where $\theta_{\scriptscriptstyle(0)}$ is the background value of the expansion scalar $\theta:=u^{\mu}{}_{;\mu}$, where $u^{\mu}$ is the four-velocity $u^{\mu}:=c^{-1}U^{\mu}$, normalized to unity, $u^{\mu}u_{\mu}=1$. A semicolon is used to denote covariant differentiation with respect to the background metric tensor $g_{{\scriptscriptstyle(0)}\mu\nu}$ (5). The equations (10c) and (10d) represent the energy density conservation law and the particle number conservation law, respectively. It is crucial to highlight that the variable $w$ in Eq. (10c) is an abbreviation for $p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, and it does not represent the equation of state. The spatial parts of the background Riemann tensor, ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}jkl}$, the Ricci curvature tensor, ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}j}$, and its contraction, ${{}^{3}\!}R_{\scriptscriptstyle(0)}$, are given by the following expressions:
@@
where ${{}^{3}\!}R_{\scriptscriptstyle(0)}$ is the spatial curvature. The value of $K$ determines the nature of the flrw universe. The universe is open for $K=-1$, flat for $K=0$, and closed for $K=+1$.

### v3

@@ IV Gauge Problem and its Solution
Einstein’s equations and conservation laws are invariant under coordinate transformations $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$. This signifies that there are no preferred coordinate systems. Consequently, the linearized Einstein equations and conservation laws are gauge-invariant, meaning they are invariant under linear, infinitesimal space-time transformations, also known as gauge transformations, which are given by
@@
where the gauge functions $\xi^{\mu}(t,\bm{x})$ are four arbitrary infinitesimal functions of time, $x^{0}=ct$ with $c$ the speed of light, and space, $\bm{x}=(x^{1},x^{2},x^{3})$, coordinates. The gauge problem of cosmology stems from the linear transformation (3) and the linearity of the linearized Einstein equations and conservation laws. This problem will now be examined.
@@
where $u^{\mu}$ is the fluid four-velocity normalized to unity, $u^{\mu}u_{\mu}=1$}, and $N^{\mu}:=nu^{\mu}$ represents the cosmological particle current four-vector, which satisfies the particle number conservation law $N^{\mu}{}_{;\mu}=0$}. A semicolon denotes covariant differentiation. Let $S_{\scriptscriptstyle(0)}(t)$ denote the quantities $\varepsilon_{\scriptscriptstyle(0)}$, $n_{\scriptscriptstyle(0)}$, and $\theta_{\scriptscriptstyle(0)}$. These quantities satisfy the background equations, that is to say, the unperturbed Einstein equations and conservation laws. Let $S_{\scriptscriptstyle(1)}(t,\bm{x})$ be their perturbed counterparts $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, which satisfy the linearized equations. As a consequence of the linearity of these equations, new solutions that have precisely the same physical content can be generated. These new solutions are given by (see Weinberg [42], Sec. 10.9 for a detailed explanation)
$S^{\prime}_{\scriptscriptstyle(1)}=S_{\scriptscriptstyle(1)}+\mathcal{L}_{\xi}S_{\scriptscriptstyle(0)}=S_{\scriptscriptstyle(1)}+\xi^{0}\dot{S}_{\scriptscriptstyle(0)},\quad S=\varepsilon,n,\theta.$ (5)
@@
In this expression, the operator $\mathcal{L}_{\xi}$ is the Lie derivative with respect to the infinitesimal four-vector $\xi^{\mu}$. An overdot denotes differentiation with respect to $x^{0}=ct$. The solutions $S_{\scriptscriptstyle(1)}$ and $S^{\prime}_{\scriptscriptstyle(1)}$ contain the so-called gauge modes which are given by $\hat{S}_{\scriptscriptstyle(1)}=\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$. This indicates that the solutions $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ to the linearized equations are gauge-dependent. Since these solutions depend on the choice of a reference frame, they are devoid of physical significance. Therefore, $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ do not represent physical phenomena. This is the gauge problem of cosmology, whereby the coordinate artifacts and the underlying physics are inextricably linked in the solution to the linearized equations. Now that the gauge problem of cosmology has been understood, a solution can be found.
@@
The physics of the perturbations is hidden in the solutions to the linearized Einstein equations and conservation laws. Consequently, quantities that describe physical phenomena must be linear combinations of gauge-dependent solutions to these equations. Since the transformation of the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ under the infinitesimal transformation (3) is identical to that described by (5), we can construct gauge-invariant quantities by combining these three gauge-dependent quantities in a way that eliminates the gauge modes $\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$. This results in three distinct sets of linear combinations. In each of these sets, precisely one gauge-invariant quantity is identically zero. Since we are only interested in energy density and particle number density perturbations, the only possible set is
@@
From these definitions, it can be inferred that the general form of a gauge-invariant quantity describing a physical phenomenon is given by
$q^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}:=q_{{\scriptscriptstyle(1)}}-\dfrac{\dot{q}_{{\scriptscriptstyle(0)}}}{\dot{\theta}_{{\scriptscriptstyle(0)}}}\theta_{{\scriptscriptstyle(1)}},$ (7)
where $q$ stands for energy density $\varepsilon$, particle number density $n$, temperature $T$, pressure $p$, or the entropy per particle $s$.
The quantities (6a) represent density perturbations and (6b) represents the perturbation to the expansion of the universe. This will be demonstrated in Sec. X on the Newtonian limit.
V Selection of a Reference Frame
@@
The quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ are defined without the use of a coordinate system. However, to derive evolution equations for these quantities, it is first necessary to choose a suitable reference frame. Since general relativity is covariant, and the quantities in question are gauge-invariant, it follows that the evolution equations for these quantities can be derived in any chosen coordinate system. The selection of an appropriate coordinate system for this problem is dependent on two criteria, which are outlined below.
@@
Firstly, to show that $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent density perturbations requires taking the Newtonian limit. In the context of general relativity, coordinate transformations are in general space-time transformations, specifically $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$. In the linear case, space-time transformations are given by (3). In Newtonian gravity, space and time are different and independent entities. Consequently, in the Newtonian limit, the relativistic space-time transformations (3) must be automatically separated into independent space and time transformations. Secondly, it would be preferable to have a coordinate system that would facilitate the derivation of the perturbation theory. It will now be demonstrated that both of these requirements can be met by employing the same coordinate system.
@@
Firstly, in Newtonian gravity, where space and time are treated as distinct and independent entities, all coordinate systems are inherently synchronous. In view of the Newtonian limit, it can be seen that a synchronous reference frame represents the optimal choice. In these coordinates, the metric tensor $g_{\mu\nu}(t,\bm{x})$ of the flrw universes is given by
@@
The scale factor of the universe is designated as $a(t)$. Given that $g_{00}=1$, coordinate time is equal to proper time. The synchronicity condition is $g_{0i}=0$, as explained in Ref. [42], Sec. 11.8 and Ref. [44], § 84. The metric tensor of the three-dimensional maximally symmetric subspaces of constant time is denoted by $\tilde{g}_{ij}$. The Killing equations $\xi_{\mu;\nu}+\xi_{\nu;\mu}=0$ and (8) demonstrate that the gauge functions $\xi^{\mu}(t,\bm{x})$ in the transformation (3) become
@@
if only transformations between synchronous coordinates are allowed (see Ref. [42], Sec. 10.9). In expressions (9), $\psi(\bm{x})$ and $\chi^{i}(\bm{x})$ are four arbitrary infinitesimal functions of the spatial coordinates.
Secondly, in synchronous coordinates the space-space components $R_{ij}$ of the four-dimensional Ricci tensor $R_{\mu\nu}$ are partitioned into two distinct parts. One part exclusively comprises the time derivatives of the spatial metric tensor $g_{ij}$ while the second part is the Ricci tensor ${{}^{3}\!}R_{ij}$ of the three-dimensional subspaces. This is demonstrated in Ref. [44], § 97. This implies that the spatial metric and Ricci tensor perturbations are both tensors. Furthermore, the spatial part of the fluid four-vector is a vector in synchronous coordinates. This allows us to apply the decomposition theorem for second-rank, symmetric three-tensors and the Helmholtz theorem to decompose the fluid vector. Using these theorems, it is shown that the linearized Einstein equations and conservation laws can be decomposed into three mutually exclusive systems: one for tensor perturbations, one for vector perturbations, and one for scalar perturbations. Next, it will be demonstrated that only scalar perturbations are associated with density perturbations. Therefore, when studying density perturbations, it is sufficient to consider the evolution equations for scalar perturbations. This system is easier to use than the original set of equations, thus facilitating the derivation of a cosmological perturbation theory. Because synchronous coordinates are also compatible with Newtonian gravity, they will be used from this point onward.
@@ VI Einstein Equations and Conservation Laws
As was concluded in the previous section, synchronous coordinates are the optimal choice for deriving a perturbation theory. This section presents the background Einstein equations and conservation laws, as well as their linearized counterparts, in this coordinate system.
The background energy density constraint equation and its linearized counterpart are expressed in a manner consistent with the contracted Bianchi identities. These identities state that the Riemann tensor satisfies the equation $R^{\mu}{}_{\nu;\mu}=\tfrac{1}{2}R_{,\nu}$, where a comma denotes ordinary differentiation. This can be expressed in an equivalent manner as the four-divergence of the Einstein tensor, i.e., $G^{\mu\nu}{}_{;\nu}=0$}. This result demonstrates that the constraint equations contain at most first-order time derivatives of the metric, which is precisely what is required for the development of a cosmological perturbation theory. For further details see Ref. [42], Secs. 6.8 and 7.5 and Ref. [44], § 92 and § 95.
@@ VI.1 Background Equations
The $G_{0i}$ constraint equations and the $G_{ij}$ dynamical equations with $i\neq j$ are satisfied identically. The $G_{ii}$ dynamical equations are identical to the time derivative of the $G_{00}$ constraint equation (Friedmann equation) (11a). Therefore, the $G_{0i}$ and $G_{ij}$ equations can be omitted. The cosmological constant $\Lambda$, the gravitational constant $G_{\text{N}}$, and the speed of light $c$ are fundamental constants.
The overdot denotes differentiation with respect to $x^{0}=ct$. The Hubble function $H$ is defined by $H:=\dot{a}/a$. For flrw universes, the Hubble function is given by $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, where $\theta_{\scriptscriptstyle(0)}$ is the background value of the expansion scalar $\theta:=u^{\mu}{}_{;\mu}$, where $u^{\mu}$ is the four-velocity $u^{\mu}:=c^{-1}U^{\mu}$, normalized to unity, $u^{\mu}u_{\mu}=1$. A semicolon denotes covariant differentiation with respect to the background metric tensor $g_{{\scriptscriptstyle(0)}\mu\nu}$ (8). The equations (11c) and (11d) represent the energy density conservation law and the particle number conservation law, respectively. The variable $w$ in Eq. (11c) is an abbreviation for $p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, and it does not represent an equation of state.
The spatial parts of the background Riemann tensor ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}jkl}$, the Ricci tensor ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}j}$, and its contraction ${{}^{3}\!}R_{\scriptscriptstyle(0)}$ are given by
@@
where ${{}^{3}\!}R_{\scriptscriptstyle(0)}$ is the spatial curvature. The value of $K$ determines the nature of the flrw universe. The universe is open for K=1K=-1, flat for $K=0$, and closed for $K=+1$.